# Problem Solving Through Human-AI Preference-Based Cooperation

### v3

@@ 3 HAI-Co2: Human-AI co-construction through preference-based search
In this section, we propose HAI-Co2, a framework for cooperative problem solving. Broadly speaking, HAI-Co2 is meant to formalize an interactive problem solving scenario, in which a human expert seeks to (co-)construct a candidate solution – such as a computer program – with the help of an AI agent. The problem solving process is conceived as a process of systematic search in a space $\mathcal{X}$ of candidate solutions, i.e., as a co-constructive process, in which candidate solutions are modified or extended step by step until a (sufficiently) good solution has been found. Therefore, we also refer to the search space $\mathcal{X}$ as the construction space. The construction space, its hierarchical organization and its topology (or neighborhood structure) are depicted in Figure 2.
@@ 3.1 Construction space and abstraction hierarchy
The construction space will typically be large, most often even (countably) infinite. For example, the construction space may consist of all computer programs in a specific programming language. Spaces of this kind cannot be specified in an explicit way. Instead, they will be defined implicitly and may even be adapted or designed on-the-fly in the course of the problem solving process. In this regard, the formal representation of candidate solutions is of major importance and will strongly influence the efficacy and efficiency of the problem solving process. Moreover, it is also clear that the representation of solutions will not be universal but rather specific to the expert domain. For example, a computer program will not be represented in the same way as a machine learning pipeline or data science workflow. It should be noted that we do not make any assumption of completeness for candidate solutions: at any stage of the search, a candidate solution ${\bm{x}}\in{\mathcal{X}}$ can be partial or incomplete (i.e., an incomplete codebase, an incomplete ML pipeline, etc.).
During problem solving, it is often useful to look at (candidate) solutions on multiple levels of abstraction. In many cases, for example, a rough draft of the solution is found in a first phase of the process, and this draft is then worked out in more detail in a second phase. More generally, one can imagine a search process that switches back and forth between different levels of abstraction whenever appropriate. Therefore, we assume the construction space ${\mathcal{X}}$ is equipped with a hierarchy of abstraction levels.
@@
In our case study on code generation presented in Section 5, we implement the construction space on three levels of abstraction, considering a Python program as a refinement of a UML diagram, which in turn is a refinement of a natural language specification. The refinement maps are implemented by suitably prompted LLMs that take a representation in a higher abstraction space as input and produce a solution representation in the lower level of abstraction as output.
@@ 3.2 Latent utility
We assume that the construction space is equipped with a latent utility function reflecting the preferences of the expert, i.e., the quality of solutions as perceived by the expert. In general, “quality” may refer to various dimensions or criteria, and different objectives might be pursued at the same time; we formalize this with a multidimensional utility function $u({\bm{x}})=(u_{1}({\bm{x}}),\ldots,u_{K}({\bm{x}}))^{\top}$ comprised of local utility functions $u_{i}$. For example, a computer program could be rated by average runtime or memory consumption. The local utility functions can be combined into a scalar utility function $U:\,{\mathcal{X}}\rightarrow\mathbb{R}$ via a suitable aggregation operator.
@@
In general, the utility (be it in the form of the multidimensional utility function $u$ or the scalar utility function $U$) is not known to the AI, nor is the expert explicitly aware of it. Rather, this utility is latent and underlies the expert’s preference feedback. From this preference feedback, the AI can learn an approximation $\hat{U}$. The AI’s goal is then to construct a solution ${\bm{x}}^{*}$ that maximizes $\hat{U}$, or which is at least close to the maximizer, while simultaneously improving the approximation quality of $\hat{U}$. The utility $U$ also induces utilities on higher levels of abstraction. For example, one way to “lift” a utility function from level ${\mathcal{X}}_{j}$ to the more abstract level ${\mathcal{X}}_{j-1}$ is via aggregation: $U({\bm{x}}^{\prime})=\alpha(\{U({\bm{x}})\,|\,{\bm{x}}\in{\mathcal{X}}_{j},f_{j}({\bm{x}})={\bm{x}}^{\prime}\})$, where $\alpha$ is an appropriate aggregation function (Grabisch et al., 2009).
In our case study, we represent the utilities as user preferences in natural language, extracted from the interaction where both the user and the AI agent can choose between presented options, provide explicit instructions, or actively edit parts of the solution.
@@ 3.3 Interaction and preference-based search
To guide the search, human and AI can communicate via natural language; e.g., the AI agent may ask the expert for feedback or explicit advice. Alternatively, the expert may actively intervene, for example by critiquing or modifying a candidate solution. A third type of interaction, particularly important in the context of HAI-Co2, is driven through preferential feedback: By informing the AI agent about the quality of candidate solutions, the expert provides hints at presumably more promising (and, likewise, less promising) regions of the construction space, and hence suggests promising “search directions” to the AI agent. To give an example, the expert can compare two competing candidate solutions with each other (e.g., whether a modification has improved a solution or made it worse) and provide this feedback (in natural language) to the AI agent for the next iteration. The AI agent utilizes the feedback to improve its approximation $\hat{U}$ of the latent utility function, which is an important element of its informational state.
In our case study, for example, we implement a preference-based search strategy that identifies promising solutions via a tournament of pairwise comparisons. Besides, we realize a search policy that refines an existing solution guided by the expert’s preferences (see Section 5 for details).
@@
  • Utility: soft/hard constraints, preference relations/predicates (i.e., what type of preferences can in general be expressed, and in which form)
@@
HAI-Co2 comprises a broad variety of human-AI interaction in natural language, as well as categorical choices and active modification of the candidate solution by the human expert. The search policy $\pi$ is designed to generate a (locally) optimal candidate solution based on the immediate as well as historical feedback, thereby adapting to the preference signals from the human expert user. The hierarchical abstraction of the search space facilitates a modular modification of the complex candidate solution. As we will see in the case study (Section 5), HAI-Co2 also allows for incorporating creative components into the generation of candidate solutions, for example through the injection of randomness in the heuristic search process or the refinement of abstract into more concrete solutions.
4 Challenges and practical issues
@@
Our characterization of HAI-Co2 implies multiple challenges that need to be addressed to realize co-construction effectively. In the following, we briefly describe these challenges with reference to the current state of research.
@@
Specification of abstraction hierarchy. Two core components of HAI-Co2 are (i) the abstraction hierarchy of the construction space and (ii) the neighborhood structure that facilitates preference-based search. A synergistic implementation of (i) and (ii) poses a non-trivial research challenge. In the domain of code generation, Le et al. (2024) propose a modular code generation approach to circumvent this challenge: they generate a chain-of-thought style intermediate description of the subtasks followed by modular codes implementing each of them. Such a hierarchical generation approach can be extended to generalized solution co-construction. However, relying on purely natural language-based intermediate representations limits the utility of the hierarchy. The choice of abstraction representation is domain-specific: UML descriptions are a suitable abstraction for code generation, but not for an AI scientific assistant. The abstraction specification must adhere to the neighborhood structure on all levels of abstraction; e.g., if two points are neighbors, their abstractions must also be neighbors. This poses additional constraints on the choice of abstraction, the choice of neighborhood structure, and the choice of refinements between two given levels of abstraction. Ideally, all levels of abstraction must be suitable for human preference expression: an “unnatural” choice of abstraction can render the co-construction tiresome for the human expert, negatively affecting productivity.
@@
Learning from active human edits. The role of the AI agent as a co-construction partner is central to HAI-Co2. This entails the possibility of active participation from the human expert and the need to learn expert preferences from such participation. Current generative AI lacks the necessary structures of the solution space, primarily represented in its input context, that could delineate the changes introduced by the expert and, subsequently, be the basis for learning from it. HAI-Co2 provides a plausible alternative to “put everything in context” that can solve this challenge, as we argue in the following. Let ${\bm{x}}\in{\mathcal{X}}_{j}$ and ${\bm{x}}^{\prime}\in{\mathcal{X}}_{j}$ be the solution before and after the edit from the human expert. The neighborhood structure imposed by HAI-Co2 on ${\mathcal{X}}_{j}$ requires learning the changes in the utility function upon moving from a candidate solution to a neighboring one. If one ensures a vector space structure formed by the utility ${\bm{u}}({\bm{x}})$, then the expert preference is equivalent to ${\bm{u}}({\bm{x}}^{\prime})-{\bm{u}}({\bm{x}})$. Even without learning to map the solutions to the utility, one can simply seek to learn the mapping from ${\bm{x}}^{\prime}-{\bm{x}}$ to expert preferences. Gao et al. (2024) previously showed that learning preferences from such changes is superior to prompting-based methods in terms of aligning LLMs to user edits. However, their experiments are focused on simpler, general-purpose natural language generation tasks. In expert-domain applications where the utility of a solution includes multiple hard constraints (e.g., executability of code) along with stylistic preferences, learning a structured representation of the utility space is essential and a challenge on its own.
@@
Communication in natural language. When humans co-construct a solution, communication in natural language plays an important role. Natural language is a powerful and at the same time succinct medium for conveying information. Given the expressivity of natural language, human and AI agent can easily communicate different options of how to improve the current solution, both at a detailed level and in more abstract terms. Similarly, preference learning is facilitated by natural language, since many preferences are easily specified in natural language. The challenge here is that the language capabilities of LLMs have advanced to an impressive level for the general domain, but this does not apply to complex expert domains (Magesh et al., 2024; Hager et al., 2024; Anand et al., 2024).
Multimodal human-AI interaction. Natural language-based interaction is not the ideal channel for all types of preferences. Categorical preference can be communicated more simply by pointing towards a better solution. Thus, we would like to incorporate multiple types of preference into HAI-Co2. Similar to deciphering the preference from natural language, different modalities and modeling approaches have their own sets of challenges and require non-trivial research efforts. For example, inferring a preference-based global ranking from pairwise comparisons can be challenging. Popular methods like the Bradley-Terry model (Bradley & Terry, 1952) have their own limitations, such as strong assumptions on the preference structure. Prior literature tackling such hurdles (Mao et al., 2018; Shah et al., 2016) paves the way for research under the umbrella of HAI-Co2. Additionally, incorporation of expert preferences across multiple modalities poses the challenge of aligning these multiple modes of feedback with each other. For example, the human expert may express the need for a security feature in a software engineering problem explicitly, or they can express it implicitly by choosing a candidate solution that includes the feature over one that does not. The AI needs to extract equivalent preference information in these two scenarios. Contemporary research in recommender systems that deal with modeling user preferences on multiple item modalities (Guo et al., 2018; Xu et al., 2021) can serve as a starting point. However, the relative complexity and nuances of preferences in the case of HAI-Co2 hinder a trivial extension of recommendation-oriented solutions.
Dynamic user preferences. Current techniques of aligning neural AI systems to human preferences, broadly referred to as RLHF (Reinforcement Learning from Human Feedback), typically involve a two-stage process: learning a reward model on preference data followed by fine-tuning a foundation model (often an LLM or a diffusion model) upon reward supervision from the reward model (Kaufmann et al., 2024). This setup is fundamentally limited to static adaptation in the regime of expert-domain co-construction; a single model of human preferences is imitated by the agent that cannot adapt to the personalized preferences of the human expert. In practice, user preferences are dynamic and evolve over time (Franklin et al., 2022). They can also be influenced by and hence be constructed during the elicitation process itself (Lichtenstein & Slovic, 2006). This is a fundamental challenge in co-construction problems, where the AI agent must adapt to the evolving preferences of the human expert. Multi-turn RLHF (Zhou et al., 2024), though it extends the context of preference-adherence to an iterative, conversational regime, cannot solve the challenge of dynamically evolving user preferences. The PAL framework, proposed by Chen et al. (2024), provides a partial solution to our problem via personalized modeling of static human preferences. Unlike traditional policy learning, HAI-Co2 motivates a reward-free exploration of the solution space (Jin et al., 2020). In-context reinforcement learning can pave the way towards handling dynamic preference signals (Yang et al., 2024; Lee et al., 2024). However, the action space in the scope of HAI-Co2 overlaps with the generation of multiple hierarchical views of the candidate solution, rendering the problem much harder than existing work on in-context RL. Prior work with LLMs showcases the possibilities of using them as in-context agents, though exploration abilities will need fine-tuning-based interventions (Krishnamurthy et al., 2024).
@@
Creativity-correctness dilemma. The specific class of co-construction problems that we seek to address requires creative generation. At the same time, in most expert-domain applications, the solution needs to fulfill objective correctness criteria. With generative models, the two requirements of creativity and correctness become counteractive. Creative generation typically emerges in highly stochastic regimes (e.g., high temperature decoding) (Wang et al., 2023a). However, increased stochasticity carries the risk of hallucination (Aithal et al., 2024). For problems with definite answers, it has already been shown that more robust reasoning can be achieved by stochastic exploration of the generation space and identifying the subset of solutions that are most consistent (Wang et al., 2023b). However, such self-consistency methods are limited to problem classes with definitive answers and cannot be readily applied to the co-construction problems that we characterize in this paper. In HAI-Co2, this can be generalized into a broader learning problem of exploration-exploitation trade-off. In the early iterations of co-construction, when the preference input from the human expert is likely to be vague, the AI may bias towards exploration of the construction space in search for a creative solution backbone. As the co-construction proceeds, the human expert fixates on the feature requirements and the AI must refrain from abrupt modifications and build on the preference model developed from the early exploration.
@@
Evaluation of co-construction techniques. Due to the personalized and dynamic preferences of the expert and the complexity and modularity of the solution, the evaluation of co-construction is a difficult challenge. We identify multiple dimensions of evaluation that need to be addressed:
@@
  • Quality of the solution should be evaluated using domain-specific measures; irrespective of the process of co-construction or human preferences, the solution must fulfill some objective criteria of correctness.
@@
  • Preference-adherence is an essential criterion of the co-construction problem; across multi-iteration co-construction, the generation should closely follow what the human expert asks of it.

### v4

@@ 3 Towards a formalization of HAI-Co2
In this section, we propose HAI-Co2, a framework for cooperative problem solving. Broadly speaking, HAI-Co2 is meant to formalize an interactive problem solving scenario, in which a human expert seeks to (co-)construct a solution – such as a computer program – with the help of an AI agent. The problem solving process is conceived as a process of systematic search in a space $\mathcal{X}$ of candidate solutions, i.e., as a co-constructive process, in which candidate solutions are modified or extended step by step until a (sufficiently) good solution has been found. Therefore, we also refer to the search space $\mathcal{X}$ as the construction space. The construction space, its hierarchical organization and its topology (or neighborhood structure) are depicted in Figure 2.
@@
As a running example, we provide the code generation use case in Figure 3 as an illustration of how different formal aspects of HAI-Co2 can be implemented (see Appendix A for more details). A second use case – in the expert domain of related work section generation – can be found in Appendix B.
Figure 3: Running example. An example instantiation of HAI-Co2 for the problem of building a double pendulum simulation. A solution is co-constructed through human-AI cooperation as follows. The interaction between the human and the AI is shown in the red box on the right, top to bottom. We define the co-construction space on three levels of hierarchy: Specification space, UML space, and Python program space. The user starts by specifying the (potentially incomplete) problem to solve. The Specification module (green rectangle) generates a pair of candidate specifications of the software to build (S1 and S2) and presents them to the user. The user expresses their preference in two manners: i) they choose one of the candidate solutions (S1) as better than the other, and ii) provide partial edits to the specification directly. The Preference decoder module (purple rectangle) extracts preference rules from the interaction context. Based on the decoded preferences, the Specification module generates a new pair of candidate specifications (only one shown for space reasons), from which the user chooses one. The UML generation module serves as a generator of refinements from specification space to UML space and generates a set of four UMLs from the specification selected. The Preference-based search module then runs a tournament-based search among these UML candidates: a pair of UMLs are compared against the specification and the decoded user preferences and one is chosen. Two finalist UMLs from the tournament are then used by the Code generation module (pink rectangle) to generate two candidate Python programs. These programs are presented to the user again for their feedback.
@@ 3.1 Construction space and abstraction hierarchy
The construction space will typically be large, most often even (countably) infinite. For example, the construction space may consist of all computer programs in a specific programming language. Spaces of this kind cannot be specified in an explicit way. Instead, they will be defined implicitly and may even be adapted or designed on-the-fly in the course of the problem solving process. In this regard, the formal representation of candidate solutions is of major importance and will strongly influence the efficacy and efficiency of HAI-Co2. Moreover, it is also clear that the representation of solutions will not be universal but rather specific to the expert domain. For example, a computer program will not be represented in the same way as a machine learning pipeline or data science workflow. It should be noted that we do not make any assumption of completeness for candidate solutions: at any stage of the search, a candidate solution ${\bm{x}}\in{\mathcal{X}}$ can be partial or incomplete (i.e., an incomplete codebase, an incomplete ML pipeline, etc.).
During problem solving, it is often useful to look at (candidate) solutions on multiple levels of abstraction. In many cases, for example, a rough draft of the solution is found in a first phase of the process, and this draft is then worked out in more detail in a second phase. More generally, one can imagine a search process that switches back and forth between different levels of abstraction whenever appropriate. Therefore, we assume that the construction space ${\mathcal{X}}$ is equipped with a hierarchy of abstraction levels.
@@
In our running example on code generation presented in Appendix A, we implement the construction space on three levels of abstraction, considering a Python program as a refinement of a UML diagram, which in turn is a refinement of a natural language specification. The refinement maps are implemented by suitably prompted LLMs that take a representation in a higher abstraction space as input and produce a solution representation in the lower level of abstraction as output. The one-to-many nature of the refinement is reflected via stochastic sampling in the LLM inference: multiple UML representations are sampled stochastically from one natural language specification. In case of related work section generation, these abstractions can be conceptualized as list of papers, semantic graph/outline of related work, and the actual related work section (the final artifact/solution, see Appendix B). Note that in either case, the abstraction hierarchy can facilitate better utilization of different expertise of the human and the AI in the co-construction process; for example, modern code-language models are great at stylizing a piece of code (variable naming, commenting, etc.), while human understanding of requirement engineering is needed to reach the optimal specification document. Similarly, an LLM paired with a search engine can possibly result in an efficient paper-finder agent in the use case of related work generation, while the human scientist defines the utility function that evaluates a candidate related work section.
@@ 3.2 Latent utility
We assume that the construction space is equipped with a latent utility function reflecting the preferences of the expert, i.e., the quality of solutions as perceived by the expert. But the utility function also reflects input the AI agent may have given on which solutions are to be preferred. Thus, the utility function embodies the current state of the co-construction of the objective that we introduced in Section 2. This also means that the utility function is not static, but changes as expert and agent refine and change their goals during the co-construction process.
What exactly do we mean by “quality” of the solution? In general, “quality” may refer to various dimensions or criteria, and different objectives might be pursued at the same time; we formalize this with a multidimensional utility function $u^{t}({\bm{x}})=(u^{t}_{1}({\bm{x}}),\ldots,u^{t}_{K}({\bm{x}}))^{\top}$ comprised of local utility functions $u^{t}_{i}$ where the time index $t$ indicates the temporal dependency and dynamic nature of the utility function. For example, a computer program could be rated by average runtime or memory consumption. The local utility functions can be combined into a scalar utility function $U^{t}:\,{\mathcal{X}}\rightarrow\mathbb{R}$ via a suitable aggregation operator.
@@
In general, the utility (be it in the form of the multidimensional utility function $u^{t}$ or the scalar utility function $U^{t}$) is not known to the expert and the AI agent, nor are they explicitly aware of it. Rather, the utility is latent and underlies the expert’s preference feedback, potentially taking into account the AI agent’s input on what makes a good solution. Based on this, the AI can learn an approximation $\hat{U}^{t}$. The AI’s goal is then to construct a solution ${\bm{x}}^{*}$ that maximizes $\hat{U}^{t}$, or which is at least close to the maximizer, while simultaneously improving the approximation quality of $\hat{U}^{t}$. The utility $U^{t}$ also induces utilities on higher levels of abstraction. For example, one way to “lift” a utility function from level ${\mathcal{X}}_{j}$ to the more abstract level ${\mathcal{X}}_{j-1}$ is via aggregation: $U^{t}({\bm{x}}^{\prime})=\alpha(\{U^{t}({\bm{x}})\,|\,{\bm{x}}\in{\mathcal{X}}_{j},f_{j}({\bm{x}})={\bm{x}}^{\prime}\})$, where $\alpha$ is an appropriate aggregation function (Grabisch et al., 2009).
In our running example, we represent the utilities as user preferences in natural language, extracted from the interaction where both the user and the AI agent can choose between presented options, provide explicit instructions, or actively edit parts of the solution.
@@ 3.3 Interaction and preference-based search
To guide the search, human and AI can communicate via natural language; e.g., the AI agent may ask the expert for feedback or explicit advice. Alternatively, the expert may actively intervene, for example by critiquing or modifying a candidate solution. A third type of interaction, particularly important in the context of HAI-Co2, is driven through preferential feedback: By informing the AI agent about the quality of candidate solutions, the expert provides hints at presumably more promising (and, likewise, less promising) regions of the construction space, and hence suggests promising “search directions” to the AI agent. To give an example, the expert can compare two competing candidate solutions with each other (e.g., whether a modification has improved a solution or made it worse) and provide this feedback (in natural language) to the AI agent for the next iteration. The AI agent utilizes the feedback to improve its approximation $\hat{U}^{t}$ of the latent utility function, which is an important element of its informational state.
In our running example, for instance, we implement a preference-based search strategy that identifies promising solutions via a tournament of pairwise comparisons. Besides, we realize a search policy that refines an existing solution guided by the expert’s preferences (see Appendix A for details). Additionally, we present a conceptual application of HAI-Co2 to another expert domain in Appendix B: the task of related work generation (Hu and Wan, 2014; Nishimura et al., 2024).
@@
  • Utility as the formalization of the co-constructed objective: soft/hard constraints, preference relations/predicates (i.e., what type of preferences can in general be expressed, and in which form)
@@
HAI-Co2 comprises a broad variety of human-AI cooperation in natural language, as well as categorical choices and active modification of the candidate solution by the human expert. The search policy $\pi$ is designed to generate a (locally) optimal candidate solution based on the immediate as well as historical feedback, thereby adapting to the preference signals from the human expert user. The hierarchical abstraction of the search space facilitates a modular modification of the complex candidate solution. As we can see in the running example (Appendix A), HAI-Co2 also allows for incorporating creative components into the generation of candidate solutions, for example, through the injection of randomness in the heuristic search process or the refinement of abstract into more concrete solutions.
4 Challenges and future research
@@
Our characterization of HAI-Co2 implies multiple challenges that need to be addressed to realize co-construction effectively. In the following, we briefly describe these challenges with reference to the current state of research. Additionally, we outline possible future research along these directions.
We start with the structural components of HAI-Co2 in terms of the construction space and interaction between the AI system and the human expert.
@@
Specification of abstraction hierarchy. Two core components of HAI-Co2 are (i) the abstraction hierarchy of the construction space and (ii) the neighborhood structure that facilitates preference-based search. A synergistic implementation of (i) and (ii) poses a non-trivial research challenge; several related criteria must be fulfilled, e.g., expressive and abstraction power of the hierarchical representations, ease of expressing human preference across different abstractions, aggregation of utility along the abstraction. In the domain of code generation, Le et al. (2024) propose a modular approach to circumvent this challenge: they generate a chain-of-thought style intermediate description of the subtasks followed by modular codes implementing each of them. Such a hierarchical generation approach can be extended to solution co-construction. However, relying on purely natural language-based intermediate representations limits the utility of the hierarchy – structured, semi-symbolic representations (e.g., UML diagrams for software) can provide better abstraction and facilitate ease of modification. Multi-agent systems adopting hierarchical workflows (Qian et al., 2024; Phan, Nguyen, and Bui, 2024) often simulate multi-level representation similar to human organizations. However, such abstractions are not formalized with the goal of co-construction. Such multi-agent systems suffer from limitations in specification and system design. We stress that future research toward designing the co-construction space needs to consider the strengths and limitations of the human expert and the AI system to best utilize the efforts of both. Simultaneously, the formal elements of the co-construction space (e.g., abstraction and refinement maps, utility aggregation, neighborhood structure) must be suitable for expressing expert preference.
Communication in natural language. When humans co-construct a solution, communication in natural language plays an important role. Natural language is a powerful and at the same time succinct medium for conveying information. Given the expressivity of natural language, human and AI agent can easily communicate different options of how to improve the current solution, both at a detailed level and in more abstract terms (Qian et al., 2024). Similarly, preference learning is facilitated by natural language, since many preferences are easily specified in natural language. The challenge here is that the language capabilities of LLMs have advanced to an impressive level for the general domain, but this does not apply to complex expert domains (Magesh et al., 2024; Hager et al., 2024; Anand et al., 2024). It is essential for the AI agent to understand what the human partner (as well as other AI agents in case of a multi-agent setup) asks and describe what it needs from them, which remains an important open problem in present human-AI interaction research (Bansal et al., 2024). In this regard, co-construction of objective and solution can serve as a point of fundamental rethinking: by formalizing a neighborhood and imposing a utility structure, it is possible to calibrate the effects of natural language-based interaction on the problem-solving process.
Preferences of the human expert constitute the most important functional component of HAI-Co2. Next, we outline the challenges related to preference extraction in complex problem solving with co-construction.
@@
Learning from active human edits. The role of the AI agent as a co-construction partner is central to HAI-Co2. This entails the possibility of active participation from the human expert and the need to learn expert preferences from such participation. Current generative AI lacks the necessary structures of the solution space, primarily represented in its input context, that could delineate the changes introduced by the expert and, subsequently, be the basis for learning from it. HAI-Co2 provides a plausible alternative to “put everything in context” that can solve this challenge, as we argue in the following. Let ${\bm{x}}\in{\mathcal{X}}_{j}$ and ${\bm{x}}^{\prime}\in{\mathcal{X}}_{j}$ be the solution before and after the edit from the human expert, respectively. The neighborhood structure imposed by HAI-Co2 on ${\mathcal{X}}_{j}$ requires learning the changes in the utility function upon moving from a candidate solution to a neighboring one. If one ensures a vector space structure formed by the utility $U^{t}({\bm{x}})$, then the expert preference is equivalent to $U^{t}({\bm{x}}^{\prime})-U^{t}({\bm{x}})$. Even without learning to map the solutions to the utility, one can simply seek to learn the mapping from ${\bm{x}}^{\prime}-{\bm{x}}$ to expert preferences under the assumption of a (locally) linear utility. Gao et al. (2024) previously showed that learning preferences from such changes is superior to prompting-based methods in terms of aligning LLMs to user edits. However, their experiments are focused on simpler, general-purpose natural language generation tasks. In expert-domain applications where the utility of a solution includes multiple hard constraints (e.g., executability of code) along with stylistic preferences, learning a structured representation of the utility space is essential and a challenge on its own.
Multimodal human-AI interaction. Natural language-based interaction is not the ideal channel for all types of preferences. Categorical preference can be communicated more simply by pointing towards a better solution. Thus, we would like to incorporate multiple types of preference into HAI-Co2. Similar to deciphering the preference from natural language, different modalities and modeling approaches have their own sets of challenges and require non-trivial research efforts. For example, inferring a preference-based global ranking from pairwise comparisons can be challenging. Popular methods like the Bradley-Terry model (Bradley and Terry, 1952) have their own limitations, such as strong assumptions on the preference structure. Prior literature tackling such hurdles (Mao, Weed, and Rigollet, 2018; Shah et al., 2016) paves the way for research under the umbrella of HAI-Co2. Additionally, incorporation of expert preferences across multiple modalities poses the challenge of aligning these multiple modes of feedback with each other. For example, the human expert may express the need for a security feature in a software engineering problem explicitly, or they can express it implicitly by choosing a candidate solution that includes the feature over one that does not. The AI needs to extract equivalent preference information in these two scenarios. Contemporary research in recommender systems that deal with modeling user preferences on multiple item modalities (Guo et al., 2018; Xu et al., 2021) can serve as a starting point. However, the relative complexity and nuances of preferences in the case of HAI-Co2 hinder a trivial extension of recommendation-oriented solutions.
Dynamic user preferences. Current techniques of aligning neural AI systems to human preferences, broadly referred to as RLHF (Reinforcement Learning from Human Feedback), typically involve a two-stage process: learning a reward model on preference data followed by fine-tuning a foundation model (often an LLM or a diffusion model) on reward supervision from the reward model (Kaufmann et al., 2023). This setup is fundamentally limited to static adaptation in the regime of expert-domain co-construction; a single model of human preferences is imitated by the agent that cannot adapt to the personalized preferences of the human expert. This is a fundamental challenge in co-construction problems, where the AI agent must adapt to the evolving preferences of the human expert. Multi-turn RLHF (Zhou et al., 2024), although it extends the context of preference-adherence to an iterative, conversational regime, cannot solve the challenge of dynamically evolving user preferences. The PAL framework, proposed by Chen et al. (2024), provides a partial solution to our problem via personalized modeling of static human preferences. Unlike traditional policy learning, HAI-Co2 motivates a reward-free exploration of the solution space (Jin et al., 2020). In-context reinforcement learning can pave the way towards handling dynamic preference signals (Yang et al., 2024; Lee et al., 2024). However, the action space in the scope of HAI-Co2 overlaps with the generation of multiple hierarchical views of the candidate solution, rendering the problem much harder than existing work on in-context RL. Prior work with LLMs showcases the possibilities of using them as in-context agents, though exploration abilities will need fine-tuning-based interventions (Krishnamurthy et al., 2024).
Guardrails for objective co-construction. Co-construction of the objective opens the door to manipulation to AI agents (Carroll et al., 2019; Hong, Levine, and Dragan, 2023); see Anthropic (2025a) for a real-world example. As we discussed in Section 2, even if there is no general solution to this problem, we see promising avenues of research in the specific context of HAI-Co2, including introducing a secondary monitoring agent that alerts the human when there is a suspicion of manipulative behavior by the primary agent and employing the methodology of AI alignment for training the agent to refrain from manipulative behavior – a methodology that is an active area of research with many challenges Casper et al. (2023), but has nevertheless been successful in reducing the risks of generative AI Ji et al. (2023); Anthropic (2025b). Another perspective is to broaden the scope of observation. As Miehling et al. (2025) argue in the context of multi-agent systems, focusing on the individual abilities of isolated agents in a multi-agent ecosystem can lead to underestimation – newer phenomena can emerge from the inter-agent and agent-environment interactions. We extend this argument to human-AI interactions and call for agent integrity research under the broad vision of co-construction of solution and objective.
Finally, the search-based methodology outlined under HAI-Co2 entails certain technical challenges inherent to the present-day AI, which we discuss in the following. Additionally, with the multi-turn feedback-driven process of problem solving, we discuss the non-triviality of evaluation as opposed to standard automatic evaluation techniques.
@@
Creativity-correctness dilemma. The specific class of co-construction problems that we seek to address requires creative generation. At the same time, in most expert-domain applications, the solution needs to fulfill objective correctness criteria. With generative models, the two requirements of creativity and correctness become counteractive. Creative generation typically emerges in highly stochastic regimes, e.g., in high temperature decoding (Wang, Liu, and Awadallah, 2023). However, increased stochasticity carries the risk of hallucination (Aithal et al., 2024). For problems with definite answers, it has already been shown that more robust reasoning can be achieved by stochastic exploration of the generation space and identifying the subset of solutions that are most consistent (Wang et al., 2023). However, such self-consistency methods are limited to problem classes with definitive answers and cannot be readily applied to the co-construction problems that we characterize in this paper. In HAI-Co2, this can be generalized into a broader learning problem of exploration-exploitation trade-off. In the early iterations of co-construction, when the preference input from the human expert is likely to be vague, the AI may bias towards exploration of the construction space in search for a creative solution backbone. As the co-construction proceeds, the human expert fixates on the feature requirements and the AI must refrain from abrupt modifications and build on the preference model developed from the early exploration.
@@
Evaluation of co-construction techniques. Due to the dynamics of the co-constructed objective and the complexity and modularity of the solution, the evaluation of co-construction is a difficult challenge. We identify multiple dimensions of evaluation that need to be addressed:
@@
  • Quality of the solution should be evaluated using domain-specific measures; irrespective of the process of co-construction (of solution and objective), the solution must fulfill some objective criteria of correctness.
@@
  • Preference-adherence is an essential criterion of the co-construction problem; across multi-iteration co-construction, the generation should be compatible with the human expert’s preference input.